The first side of a triangle is four less than twice the second side. The third side is 5 more than the second side. If the perimeter of the triangle is 217, what are the dimensions of the sides?
Answer by JulietG(1812)

A + B + C = 217
A = 2B - 4
C = B + 5
(2B-4) + B + (B+5) = 217
4B + 1 = 217
4B = 216
B = 54
If B is 54, then c is 59, and A is 104
54 + 59 + 104 = 217